OpenStudy (jmansedg):

William got $150 for making all A’s on his report card. He wants to spend it on CDs and DVDs. The local video store is having a sale; all CDs are $10 and all DVDs are $15. Which of the following is a possible combination of CDs and DVDs? A. 2 CDs and 9 DVDs B. 7 CDs and 5 DVDs C. 8 CDs and 6 DVDs D. 11 CDs and 3 DVDs

OpenStudy (jmansedg):

this is for a final i want to get a b in so get it right

OpenStudy (anonymous):

Ok so A) you can cross out because he has $150 and 9*15= 135 and 2*10= 20 and 20+135= 155 so he wouldn't be able to buy 2cd and 9 dvds

OpenStudy (jmansedg):

c?

OpenStudy (isaidavila):

Easy just multiply the quantity of CD's and DVD's for its respective cost. And sum up. If the answer países 120 then its wrong , if not its right

OpenStudy (anonymous):

the answer is b 7*10=70 , 5*15=75 and 75+70=145 so he would be able to 7 cds and 5 dvds

OpenStudy (jmansedg):

ok messg me mate please

OpenStudy (anonymous):

ok